OpenStudy (anonymous):

A rectangular prism has length x + 2, width x + 1, height 4, and volume 24. Find the width. a. 3 b. 2 c. 12 d. 16

OpenStudy (xmissalycatx):

I would say it is A

OpenStudy (radar):

I would say @xMissAlyCatx is guessing....You closed this question to quickly. You should of checked out that answer "A"............Lets say that it is right width = 3 So we then know x + 1 is 3 so x must be 2. That be the case length is 2 + 2 = 4 height is given as 4 so volume is lwh or 4*3*4 which equals 48 Oh Oh Volume is 24 somebody goofed!

OpenStudy (radar):

But do note that @xMissAlyCatx wrote "I would say it is A" not that it is actually A.

OpenStudy (xmissalycatx):

I didn't guess I put it into a formula but that's why I said "I would say" .. because I wasn't 100% sure. I'm glad that you corrected my mistake. I was just giving a thought.

OpenStudy (radar):

Very good @xMissAlyCatx , it is not your fault that the question was closed without verifying the suggested answer.

OpenStudy (radar):

Here is a solution, first we are going to solve for x, then solve for w. Volume V is equal to length times width times height, or V=lwh, we now substitute the given info in that formula: 24 = (x + 2)*(x + 1)*4 divide both sides of the equal sign by 4 giving you 6 = (x + 2)*(x + 1) mutiplying out the right hand side gives you \[6=x ^{2}+3x+3\] Now subtract 6 from both sides of equal sign\[0=x ^{2}+3x-4\] Rearrange in normal quadratic fashion getting:\[x ^{2}+3x-4=0\] This factors nicely (x+4)(x-1)=0 Gives us two values for x, we discard the negative solution and keep the x = 1. Now solve for w......

OpenStudy (radar):

w = x + 1 = 1 + 1 = 2 but now verify V=lwh= (x+2)(x + 1)4=3*2*4 = 24 24=24

OpenStudy (radar):

Take time to learn.

OpenStudy (anonymous):

thank you for helping me

OpenStudy (radar):

You are welcom, and good luck with your studies.